Lane Company manufactures a single product that requires a great deal of hand labor. Overhead cost is applied on the basis of1 Reqz Reg 1 Reg 2 Req 3A Req3A Req 3B Req 3B Rega Reg 4 Compute the predetermined overhead rate for the year. Break the rateReg 1 Reg 2 Req 3A Req 3B Reg 4 Prepare a standard cost card for the companys product. (Round your answers to 2 decimal placReg 1 Reg 2 Req 3A Req 3B Reg 4 Compute the standard direct labor-hours allowed for the years production. Standard direct laReg 1 Reg 2 Req 3A Req 3B Reg 4 Complete the following Manufacturing Overhead T-account for the year. Manufacturing OverheadReg 1 Reg 2 Req 3A Req 3B Req 4 Determine the reason for any underapplied or overapplied overhead for the year by computing t
Answer #1
1)
predetermined overhead rate 15.2
variable rate 5.6
fixed rate (2,880,000/300000) 9.6
2)
Direct materials 4 pounds at 12 per pound 48
Direct labor 1.5 DLH'S 13.8 per DLH 20.7
Variable overhead 1.5 DLH'S at 5.6 per DLH 8.4
Fixed overhead 1.5 DLH'Sat 9.6 per DLH 14.4
Standard cost per unit 91.5
3)
Standard direct labor hours allowed for the years production=
Actual units produced *Direct labor hr per unit
240000*1.5
360000
3b)
Manufacturing overhead T-Account
Actual costs 4368000 Applied costs (360000*15.2) 5472000
overhead overapplied 1104000
4)
Variable overhead rate variance = (AR- Std rate)*AH
                                                           =(1,248,000-390000*5.60)
936000 F
Variable overhead efficiency variance = (AH-SH)*SR
                                                                     = (390,000-360000)*5.6
168000 U
Fixed overhead budget variance = actual overhead - budgeted oveerhead
3,120,000-2,880,000
240000 U
fixed overhead volume variance= actual output*fixed overhead rate per unit-budgeted overhead
240,000*14.4-2,880,000
576,000 F